### v1

# Automated Privacy Information Annotation in Large Language Model Interactions

Abstract
Users interacting with large language models (LLMs) under their real identifiers often unknowingly risk disclosing private information. Automatically notifying users whether their queries leak privacy and which phrases leak what private information has therefore become a practical need. Existing privacy detection methods, however, were designed for different objectives and application scenarios, typically tagging personally identifiable information (PII) in anonymous content. In this work, to support the development and evaluation of privacy detection models for LLM interactions that are deployable on local user devices, we construct a large-scale multilingual dataset with 249K user queries and 154K annotated privacy phrases. In particular, we build an automated privacy annotation pipeline with cloud-based strong LLMs to automatically extract privacy phrases from dialogue datasets and annotate leaked information. We also design evaluation metrics at the levels of privacy leakage, extracted privacy phrase, and privacy information. We further establish baseline methods using light-weight LLMs with both tuning-free and tuning-based methods, and report a comprehensive evaluation of their performance. Evaluation results reveal a gap between current performance and the requirements of real-world LLM applications, motivating future research into more effective local privacy detection methods grounded in our dataset [^1] The code are public at https://github.com/Nidryen-zh/PrivacyAnnotation and the dataset can be found at https://huggingface.co/datasets/Nidhogg-zh/Interaction__\__Dialogue__\__with__\__Privacy..
1 Introduction
Figure 1: An example of privacy exposure.
Large language models (LLMs) have demonstrated immense potential in revolutionizing user interactions with artificial intelligence. However, during cloud-based LLM interactions, users often unintentionally expose privacy to service providers privacyconcern ; survey ; mope ; assessingprivacy ; LPLM ; beyond ; decoding . A case study of the ShareGPT corpus sharegpt reveals that nearly one-third of dialogues contain privacy leaks. One such instance of exposing the user’s home address is illustrated in Figure 1. Even more concerning, LLM service providers typically require access to users’ or organizations’ identity information, exposing unique identifiers before any interaction occurs. This real-identity interaction setting significantly increases the risk of privacy information disclosure. Therefore, proactively alerting users to potential privacy exposure during interactions with LLMs has become an urgent and practical need. When a user launches a query, whether the query contains private information or not should be informed. If so, the specific phrases in the query responsible for the privacy leakage need to be annotated. To further assist users who may still be uncertain, a concise explanation of what private information will be exposed by each annotated phrase should be provided. These notifications can help users quickly understand the nature of privacy leakage and take appropriate action, such as rewriting the query.
Previous research across different domains, including medical privacy medical1 ; medical2 ; medical3 , code privacy finding , and employment information Multilingual ; occupational , primarily focuses on detecting privacy risks of revealing individuals’ identities in anonymous online content or documents selfdisclosures ; entity2 ; synthetic . However, these scenarios differ significantly from privacy detection in LLM interactions. By authenticating identifiers, users pose queries to LLMs with their real names. It is important to detect privacy information related to the current users from these queries. If users disclose private information about unrelated third parties, such information will not be extracted since it does not constitute a harmful breach of privacy for the current user. Additionally, many existing studies are constrained to a limited set of topics or categories ner1 ; anonymisation ; categorylimit1 ; categorylimit2 . In contrast, privacy information in LLM interactions is much more diverse and cannot easily be classified into a limited set of predefined types.
To enable privacy detection for real-time user queries, it is desired to deploy an efficient model on the user side for automatic local detection before sending queries to cloud-based LLMs. The development and evaluation of such a model require large-scale datasets to ensure good generalization ability. However, there does not exist privacy annotated datasets for user interaction dialogues. Existing datasets were manually labeled and limited to particular domains or categories, with only a few thousand entries tab ; selfdisclosures ; i2b2 ; cegs ; codealltag . To address this, we design an automated scalable privacy annotation pipeline to extract privacy phrases and summarize the corresponding privacy information for user dialogues. With the assistance of GPT-4, our pipeline operates through four systematic steps: (1) privacy leakage classification, which filters the input to determine whether the user’s query leaks privacy; (2) extensive privacy categories extraction, which identifies privacy categories to guide the extraction of phrases; (3) privacy phrase extraction, where privacy phrases are extracted from queries and filtered; and (4) privacy information annotation, which summarizes details for each phrase. Based on the automated pipeline, we extract privacy phrases and annotate privacy information for dialogue datasets in different languages, containing 97,659 user queries, 85,320 phrases and corresponding information for English portion and 151,988 user queries, 68,910 phrases and information for Chinese portion. We also conduct human evaluation, demonstrating that the results obtained through our pipeline are highly consistent with those manually annotated. Compared to directly prompting LLM to extract privacy phrases ghost , our pipeline produces more accurate and diverse annotations.
The constructed dataset can serve not only as a benchmark for evaluation but also as a valuable source of training data for local privacy detection methods. For evaluation, we propose new metrics to privacy detection performance at the query, phrase, and information levels. We also set up tuning-free baselines, including prompting and in-context learning with locally deployable, light-weight LLMs, as well as supervised fine-tuning baselines. Benchmark results show that the fine-tuned 1B model with our dataset outperforms the directly prompted 72B model.
We summarize the key contributions as follows: (1) We identify a new application requirement of privacy detection for real-name user interactions with LLMs; (2) we build the automated privacy annotation pipeline to classify privacy leakage, extract privacy phrase and information, resulting in the first large-scale dataset for privacy detection in real-name interaction dialogues. We also design new metrics for benchmark testing; and (3) we report the performance of local privacy detection baselines, which, despite their progress, still fall short of meeting practical requirements. This highlights the need for further research into more effective methods grounded in our dataset. Specifically, the best-performing fine-tuned model achieves 87.6% accuracy in privacy leakage classification, a 74.3% F1 score in privacy phrase extraction, and a 44.7% F1 score in privacy information summarization.
2 Related Work
Privacy Risks of LLMs. Recent work constructs the corpus with online personal information or user interaction history privacysurvey2 ; finetuneingleak to train LLMs characterllm ; exploitingpersona , and enhance inference ability learningretrieval ; hallucination , resulting in the issue of memorization memorization ; privacysurvey1 ; ProPILE ; extractingtraining , which recalls the personal sequences from LLM corpus. Existing works solve this problem on the LLM service provider side, including data pre-processing approaches (e.g., sanitizing training data sanitizing and reducing the frequency of private data appearing Deduplicating1 ; Deduplicating2 ), and privacy-preserving training methods (e.g., differentially private training differential1 ; differential2 ; differential3 ). In contrast, we takes a different angle for privacy risks of LLMs by locally detecting privacy in the users’ interactions to prevent private information uploaded to LLM service providers, solving the problem at the root of the the user side.
Privacy Detection. Most previous work focuses on privacy detection for specific documents or context shared by users on the Internet semantics ; privacypreserving , detecting privacy that can identify an individual for anonymization entity1 ; entity2 ; synthetic . However, privacy in user interaction dialogues is fundamentally different from these scenarios. During the interaction, users have already authenticated their identifiers. The target application scenario of this work is to detect any newly leaked private data related to the user after a query is made, extending beyond just identifiable information. Existing approaches are also typically constrained to specific topics or categories anonymisation ; personalinformation , such as medical privacy medical1 ; medical2 ; medical3 ; medical4 , personal employment information Multilingual ; occupational , online comments selfdisclosures , and code privacy finding . In contrast, the types of privacy information in user interaction is not limited. Furthermore, most existing privacy detection models are based on discriminative models, treating the detection process as a sequence tagging task selfdisclosures ; ner1 ; entity3 . They can not generate specific privacy information to remind user. MemoAnalyzer ghost directly prompt generative LLMs to identify open privacy information without predefining categories, but the detection performance remains suboptimal.
Privacy Datasets. Existing datasets for privacy detection are limited to specific domains and there is no large-scale dataset for real-name user interaction dialogues. For instance, PrivacyBot privacybot provides a sentence-level privacy classification dataset, while I2B2/UTHealth i2b2 and CEGS N-GRID cegs focus on privacy entity recognition in the personal health domain. CodE ALLTAG codealltag targets privacy entity recognition in German emails, TAB tab annotates privacy entities in court judgments, and JobStack jobstack specializes in privacy entity annotation in job postings. Additionally, these datasets face challenges of data volume and they are manually labeled. In contrast, our dataset focuses on LLM interactions and is annotated automatically, resulting in significantly larger scale compared to existing datasets and encompassing a broader range of privacy. Furthermore, it offers detailed information for each extracted privacy phrase. The comparison details are provided in Appendix D.
3 Problem Formulation
Privacy in LLM Interactions. We focus on the scenario where the user interacts with LLMs by submitting real-name text queries. The LLM service providers are honest-but-curious, which implies that they may attempt to infer privacy information from queries and do not actively launch attacks. Different from PII, which refers to data that can identify an individual in an anonymous text, we define privacy information in LLM interactions as the service provider’s knowledge about user-related private data or personal details, before and after the user poses a query with real unique identifier. The definition of privacy information leakage is consistent with the principle of differential privacy, which considers the difference of the query output with and without each individual data record. The privacy information in real-name user dialogues is regarded as the information, if exposed, will harm the personal, social, or professional aspects of the user who poses the query or the user’s related
Figure 2: Automated privacy annotation pipeline with LLM API and privacy detection module with light-weight LLM.
people, such as family, friends, or close associates. Typical privacy information includes, but is not limited to, preferences, intentions, and opinions. We emphasize that the privacy information in LLM interactions should be related to the user who pose the query with direct evidences. For example, as shown in Figure 1, the statement “Shanghai is also where David works.” leaks PII about David, but it does not leaks the privacy information about the user, since there is no direct evidence indicating David is related to the user.
Automated Privacy Annotation Requirement. As shown in Figure 2, it is necessary to first build an automated privacy annotation pipeline based on a cloud-side LLM API to annotate dialogue datasets with privacy phrases and specific privacy information. Such a dataset can serve as an evaluation benchmark for downstream applications and provide training data for optimization. Local privacy detection methods with light-weight LLM are then designed to detect privacy from user real-time queries before they are sent to service providers and they should be deployed locally on the user’s device or be a request proxy built by the user’s organization. Formally, we let $D$ denote the user dialogue and $q$ denote a query from $D$. The goal is to extract the privacy phrase $\mathcal{P}_{q}=\{p_{1},...,p_{m}\}$ and summarize specific privacy information $\mathcal{I}_{q}=\{I_{1},...,I_{m}\}$, where $m$ is the number of the extracted phrases, $p_{j}$ is defined as a segment of $q$ that discloses private or sensitive information about the user, and $I_{j}$ is a statement summarizing the specific details revealed by $p_{j}$.
4 Automated Privacy Annotation Pipeline for Dialogue Datasets
We design an automated four-step pipeline based on strong LLM API to extract privacy phrase and annotate privacy information for dialogue datasets. The comparison between the results of our automated pipeline and manual annotation demonstrates a high degree of consistency.
4.1 Pipeline Rationale
A straightforward approach to extract privacy information is directly prompting GPT-4. However, direct prompting may force GPT-4 to extract information from samples that do not contain any privacy leaks, which results in a decrease in precision. Furthermore, the original understanding of privacy by GPT-4 revolves around the idea of whether a privacy phrase can reveal a user’s identifiable information. While this is intuitive for identifying PII, it does not align with our intended focus in Section 3, which aims to detect privacy information in real-name user interactions. We also observe that the results of direct prompting are limited to a few common types of privacy, such as location, name, and gender, with less accuracy in detecting more nuanced aspects like user status or intent.
To address these limitations, we developed the privacy annotation pipeline, as shown in Figure 3. (1) To avoid privacy extraction for user dialogue that do not involve privacy leakage, the first step is a filter, classifying dialogues into two categories: leaked and non-leaked. (2) The second step aims for comprehensive coverage of privacy categories. We independently extract privacy categories for each sample and take the union of all categories as the extensive set of privacy categories at the dataset level. (3) Then, with the assistant of the extensive category set, we extract privacy phrase for each sample based on the whole context of queries and filter the results based on the rules of being user-related and having clear reference. (4) The last step is to annotate privacy information for each extracted phrase. We also compare the performance of direct prompting with our automated pipeline in Section 4.3. All associated prompts can be found in Appendix A.
4.2 Pipeline Design
Privacy Leakage or Not Classification. We explicitly address the scenario where user interacts with identifier exposure and prompt GPT-4 to assess whether the user has disclosed any new privacy information during the interaction. We further clarify that privacy information must be linked to the user or individuals associated with them, and it encompasses various aspects such as opinions, preferences, and personal details. To ensure more reliable inferences, we not only ask GPT-4 to generate a simple yes or no judgment, but also request it to explain the reasoning behind its decision. Additionally, we provide a set of manually annotated samples as reference to guide the model.
Figure 3: Overview of our automated pipeline to extract privacy phrase and annotate privacy information over dialogue datasets: ➊ determining whether the input leaks privacy; ➋ extracting privacy categories from user queries that leaks privacy; ➌ extracting privacy phrases from user queries with the help of privacy categories; and ➍ annotating privacy information for each phrase.
Extensive Privacy Categories Extraction. We perform basic extraction for each sample labeled as leaking privacy. We prompt GPT-4 to directly identify privacy phrases and the corresponding categories. We also provide an example to guide GPT-4. These results are then merged to create an extensive set of privacy categories. However, since extraction for each sample is independent, some categories may be expressed differently but hold the same meaning. For example, we may encounter both “hobby” and “interests and hobbies,” which should be treated as equivalent. To address this, we use GPT-4 to identify such duplicates and assist in the merging process.
Specifically, as shown in Figure 3, we partition the shuffled privacy categories into smaller, more manageable blocks, and deduplicate categories within each block independently. The deduplicated results from all blocks are then merged. This process is repeated iteratively until no further deduplication occurs, meaning the number of categories remains unchanged before and after the deduplication step. To reduce randomness, we run the algorithm twice and perform a final deduplication by prompting GPT-4 to deduplicate across all categories at once. In practice, we find that using only one-tenth of the data for category extraction yields results that are nearly saturated in terms of diversity.
Privacy Phrase Extraction. For each sample that contains privacy leakage, based on the expensive set of privacy categories, we prompt GPT-4 to extract phrases that explicitly reveal privacy information about the user. By inserting categories into the prompts, we ask GPT-4 to match the phrases and the categories. The extracted phrase must be directly copied from the query, and the relationship between the phrase and the user should be evident, not speculative. For example, phrases should not be extracted with the reason like “the user may be interested in something.” Additionally, any phrases falling outside the given privacy categories, which may indicate an wrong extraction, should be excluded. To enhance the precision of the extracted phrases, we apply two filtering rules: (1) the privacy phrase must be directly associated with the user or their close associates, and (2) the phrase must have an explicit reference, rather than be a general phrase such as “go a place” or “have ideas”. We use GPT-4 to evaluate these criteria and remove any phrases that do not meet the rules. The results of experiments on filtering using these rules are shown in Table 1.
In practice, samples misclassified as leaking privacy will be corrected if no phrases are extracted. Furthermore, including too many categories in a single input can reduce both the precision and the recall of phrases. To mitigate this, we divide the categories into smaller blocks and make multiple requests to GPT-4 for each sample, then merge the results. This approach may result in the extraction of synonymous phrases across different requests, such as “plan to play football” and “play football.” Therefore, we prompt GPT-4 to deduplicate the results, retaining the more concise and clear phrase.
Privacy Information Annotation. For each extracted phrase, we prompt GPT-4 to annotate the specific privacy information it reveals, based on the context of whole text. For example, as shown in Figure 1, the privacy information associated with the phrase “London” is “The user lived in London when he was a child.” We also provide several manually constructed examples for annotation.
4.3 Pipeline over Dialogue Datasets
Based on the automated annotation pipeline, we create a large dataset comprising over 249K queries from 33K dialogues with 154K privacy phrases and corresponding information in different languages.
Raw Datasets. We use a diverse set of source corpora, including one English dialogue dataset, ShareGPT sharegpt , and three Chinese dialogue datasets, CrossWOZ crosswoz , DuConv duconv , and LCCC-base lccc . The large scale and varied scope of these datasets make them well-suited as our data source. Specifically, ShareGPT naturally consists of human-to-machine interactions, while the tree Chinese datasets consist of human-to-human dialogues. To adapt the human-to-human dialogues for our task, we designate the first speaker as the target user for privacy information extraction and the second speaker as the assistant. Detailed introduction of source corpora is provided in Appendix B.
Figure 4: Top 20 privacy categories in ShareGPT.
Annotated Datasets. After each step of the automated annotation pipeline, we have: (1) classifying privacy leakage obtains over 41K samples with privacy leakage and 56K without privacy leakage in the English portion of dataset, and 64K samples with privacy leakage and 88K without privacy leakage in the Chinese portion; (2) extracting privacy categories from 2,000 English samples and 5,000 Chinese samples results in 325 categories for English and 149 for Chinese; (3) extracting privacy phrases and annotating information obtain 97,695 English samples and 151,988 Chinese samples, of which 32,814 English samples and 43,255 Chinese samples contain privacy, corresponding to 85,320 privacy phrases and information in English and 68,910 in Chinese. Appendix C shows a data example and the usage details of LLM API are presented in Appendix L.
Figure 5: Word cloud of annotated privacy information.
Dataset Analysis. Taking the English dataset as a case study, Figure 4 shows the relative frequency of top 20 different privacy categories. We can find that in open-domain interaction dialogues, information such as users’ status, opinions, preferences, and other potentially sensitive details is more likely to be inadvertently revealed, whereas categories like names appear less frequently.
Figure 5 presents a word cloud of the identified privacy information in English dataset. We can observe that the word cloud reveals a high frequency of terms associated with the user’s “involvement” in various activities, as well as their personal interests. These results indicates that the privacy information in user interactions is significantly different from PII in specific documents or content on the web.
Comparison with Human Annotation. To evaluate the privacy leakage classification step, we have 8 annotators classify 200 samples from ShareGPT, labeling them as either leak or non-leak. The detailed guidelines to ensure the objectivity and consistency of the annotation results are provided in Appendix F. The results show an accuracy of 0.95, a recall score of 0.94, and a precision score of 0.97. These metrics demonstrate that results of the first step closely align with the manual annotations.
Table 1: Human evaluation results of phrase extraction, prompting GPT-4 vs. our pipeline.
[…]
Directly prompting GPT-4 68.08 90.99 77.89 +00.00
Phrase extraction in pipeline 98.67 82.90 90.10 +12.21
      + Filtering rules 96.30 90.17 93.13 +15.24
We also conduct a human evaluation for the privacy phrase extraction step, as shown in Table 1. Specifically, we manually extract privacy phrases for 100 samples to serve as the ground truth, resulting in 222 privacy phrases. Using the evaluation method described in Section 4.4, we found that the recall score for phrases directly extracted by GPT-4 with prompts ghost was only 68.08%, meaning that roughly one-third of privacy phrases were missed. In comparison to this baseline, our pipeline, by incorporating category information as additional input, shows a decrease in precision but significantly increases the number of extracted phrases, boosting the recall score by 30.59%. As a result, the F1 score improves by 12.21%. Furthermore, applying filtering rules further enhances the F1 score by an additional 3.03%. These results suggest that the phrase extraction step aligns well with manual annotations, highlighting its effectiveness and consistency.
4.4 Evaluation Metrics
Dialogue datasets with privacy annotation provide ground truth of privacy leakage, privacy phrases, and corresponding information for local privacy detection methods. We design a hierarchical set of metrics spanning three levels, each aligned with tasks of increasing granularity in privacy detection.
Privacy Query-Level Metric for Leakage Classification Task. Given an user’s query, the objective of privacy leakage classification task is to determine whether the query reveals any privacy of the user or user-related individuals. The automatically annotated data can be divided into either “leaked” or “non-leaked,” based on whether there are privacy phrases extracted. For such query-level classification task, we adopt accuracy to measure the performance of different local privacy detection methods.
Privacy Phrase-Level Metrics for Phrase Extraction Task. Given a private query, the objective of privacy phrase extraction task is to extract phrases based on the context of query that results in user’s privacy leakage. For each data sample, we focus on both missed and incorrect extractions. For missed extractions, we define the recall score of sample $q$ as the ratio of the intersection between the predicted phrases and the ground truth, normalized by the total number of ground truth phrases. For incorrect extractions, we define the precision score of $q$ as the ratio of the intersection between the predicted phrases and the ground truth, normalized by the total number of predicted phrases. The overall recall and precision scores are calculated as the average of these scores across all samples.
Specifically, we let $\mathcal{Q}_{r}$ denote the set of user queries where the extracted phrase is non-empty in the ground truth and $\mathcal{Q}_{p}$ denote the set of user queries where the extracted phrase is non-empty in the prediction results. We let $\hat{\mathcal{P}}_{q}$ and $\mathcal{P}_{q}$ denote the set of the predicted phrases and the ground truth of the query $q$, respectively. Then, the recall score $\text{R}_{\mathcal{P}}$ and precision score $\text{P}_{\mathcal{P}}$ is defined as
$\text{R}_{\mathcal{P}}=\frac{1}{|\mathcal{Q}_{r}|}\sum_{q\in\mathcal{Q}_{r}}\frac{|\hat{\mathcal{P}}_{q}\cap\mathcal{P}_{q}|}{|\mathcal{P}_{q}|},\ \ \text{P}_{\mathcal{P}}=\frac{1}{|\mathcal{Q}_{p}|}\sum_{q\in\mathcal{Q}_{p}}\frac{|\hat{\mathcal{P}}_{q}\cap\mathcal{P}_{q}|}{|\hat{\mathcal{P}}_{q}|},$ (1)
where ||$\cdot$|| ⋅ | is the size of the set. In practice, the extracted phrases may refer to the same entity but have different expressions. To address this issue, our phrase matching strategy goes beyond simply checking whether two phrases are identical. It also considers two phrases a match if one is part of the other or if their Rouge-L score exceeds 0.5 privacyimplications .
Privacy Information-Level Metrics for Information Summarization Task. Given a query with privacy phrase, the objective of privacy information summarization task is to summarize specific information that be leaked from the query. In general, we use Rouge-L to evaluate the quality of generated privacy information. Similarly, we define the miss and incorrect extractions of the privacy information. We let ^$\hat{\mathcal{I}}_{q}$ and $\mathcal{I}_{q}$ denote the set of the predicted information and the ground truth of the query $q$, respectively. For one privacy information ${I}_{j}\in{\mathcal{I}}_{q}$, the recall score is defined as $f_{\text{R}}({I}_{j},\hat{\mathcal{I}}_{q})=\max_{\hat{I}_{k}\in\hat{\mathcal{I}}_{q}}\text{{Rouge}}({I}_{j},\hat{I}_{k}),$ where, Rouge is the function that calculate the Rouge-L score for two sentences. The recall score of ${\mathcal{I}}_{q}$ is defined as the average recall score of each piece of information it contains. For $\hat{I}_{j}\in\hat{\mathcal{I}}_{q}$, the precision score is defined as $f_{\text{P}}(\hat{I}_{j},\mathcal{I}_{q})=\max_{I_{k}\in\mathcal{I}_{q}}\text{{Rouge}}(\hat{I}_{j},I_{k}).$ The precision score of ^$\hat{\mathcal{I}}_{q}$ is defined as the average precision score of each piece of information it contains. Then we can define the recall score $\text{R}_{\mathcal{I}}$ and the precision score $\text{P}_{\mathcal{I}}$ for all queries as follow,
$\text{R}_{\mathcal{I}}=\frac{1}{|\mathcal{Q}_{r}|}\sum_{q\in\mathcal{Q}_{r}}\frac{1}{|\mathcal{I}_{D}|}{\sum_{I_{j}\in\mathcal{I}_{q}}f_{\text{R}}({I}_{j},\hat{\mathcal{I}}_{q})},\ \ \text{P}_{\mathcal{I}}=\frac{1}{|\mathcal{Q}_{p}|}\sum_{D\in\mathcal{Q}_{p}}\frac{1}{|\hat{\mathcal{I}}_{q}|}{\sum_{\hat{I}_{j}\in\hat{\mathcal{I}}_{q}}f_{\text{P}}(\hat{I}_{j},{\mathcal{I}}_{q})}.$ (2)
Intuitively, $\text{R}_{\mathcal{I}}$ measures the completeness of the extracted privacy information, while $\text{P}_{\mathcal{I}}$ evaluates the accuracy of the privacy information.
After getting the recall and precision score, we can also define the F1 score for both phrase-level and information-level as follow,
[…]
5 Evaluation for Local Privacy Detection Methods
Table 2: Accuracy of different baselines with query-level metric (%).
Method Model English Chinese
Acc Acc
ZG Llama-3.2-1B $61.18_{\pm 0.49}^{*}$ $58.08_{\pm 0.49}^{*}$
Llama-3.2-3B $65.07_{\pm 0.48}^{*}$ $66.97_{\pm 0.47}$
Qwen2.5-1.5B $59.37_{\pm 0.49}^{*}$ $57.66_{\pm 0.49}^{*}$
Qwen2.5-3B $65.77_{\pm 0.47}^{*}$ $65.25_{\pm 0.48}^{*}$
Qwen2.5-7B $68.98_{\pm 0.46}^{*}$ $69.67_{\pm 0.46}^{*}$
Qwen2.5-72B $73.45_{\pm 0.44}$ $75.90_{\pm 0.43}$
ICL Llama-3.2-1B $64.01_{\pm 0.48}^{*}$ $58.32_{\pm 0.49}^{*}$
Llama-3.2-3B $60.24_{\pm 0.49}^{*}$ $61.61_{\pm 0.49}^{*}$
Qwen2.5-1.5B $63.60_{\pm 0.48}^{*}$ $63.49_{\pm 0.48}^{*}$
Qwen2.5-3B $66.42_{\pm 0.47}^{*}$ $61.56_{\pm 0.49}^{*}$
Qwen2.5-7B $68.28_{\pm 0.47}^{*}$ $72.21_{\pm 0.45}^{*}$
Qwen2.5-72B $71.17_{\pm 0.45}$ $76.83_{\pm 0.42}$
SFT Llama-3.2-1B $78.69_{\pm 0.41}$ $80.99_{\pm 0.39}$
Llama-3.2-3B $81.63_{\pm 0.39}$ $81.32_{\pm 0.39}$
Qwen2.5-1.5B $82.46_{\pm 0.38}$ $86.47_{\pm 0.34}$
Qwen2.5-3B $82.59_{\pm 0.38}$ $87.49_{\pm 0.33}$
Qwen2.5-7B $\textbf{83.57}_{\pm 0.37}$ $\textbf{87.56}_{\pm 0.33}$
Table 3: Performance of different privacy detection baselines with phrase-level metrics (%).
Method Model English Chinese
[…]
ZG Llama-3.2-1B $29.67_{\pm 0.43}^{*}$ $27.21_{\pm 0.38}^{*}$ 28.39 $22.07_{\pm 0.40}^{*}$ $26.08_{\pm 0.43}^{*}$ 23.91
Llama-3.2-3B $37.51_{\pm 0.42}^{*}$ $27.86_{\pm 0.42}^{*}$ 31.97 $52.35_{\pm 0.44}^{*}$ $26.20_{\pm 0.43}^{*}$ 34.92
Qwen2.5-1.5B $36.36_{\pm 0.42}^{*}$ $29.85_{\pm 0.43}^{*}$ 32.79 $26.73_{\pm 0.43}^{*}$ $25.47_{\pm 0.41}^{*}$ 26.08
Qwen2.5-3B $24.51_{\pm 0.38}^{*}$ $34.09_{\pm 0.45}^{*}$ 28.52 $20.46_{\pm 0.38}^{*}$ $35.07_{\pm 0.45}^{*}$ 25.84
Qwen2.5-7B $42.83_{\pm 0.43}^{*}$ $42.27_{\pm 0.46}^{*}$ 42.55 $42.65_{\pm 0.46}^{*}$ $44.74_{\pm 0.48}^{*}$ 43.67
Qwen2.5-72B $61.81_{\pm 0.44}$ $44.00_{\pm 0.43}$ 51.41 $72.37_{\pm 0.41}$ $50.14_{\pm 0.46}$ 59.24
ICL Llama-3.2-1B $10.04_{\pm 0.28}^{*}$ $23.93_{\pm 0.39}^{*}$ 14.15 $20.37_{\pm 0.38}^{*}$ $23.50_{\pm 0.40}^{*}$ 21.83
Llama-3.2-3B $44.04_{\pm 0.42}^{*}$ $28.73_{\pm 0.43}^{*}$ 34.77 $60.26_{\pm 0.45}^{*}$ $27.96_{\pm 0.42}^{*}$ 38.20
Qwen2.5-1.5B $38.85_{\pm 0.42}^{*}$ $32.78_{\pm 0.44}^{*}$ 35.56 $39.81_{\pm 0.47}^{*}$ $25.99_{\pm 0.40}^{*}$ 31.45
Qwen2.5-3B $51.37_{\pm 0.41}^{*}$ $30.91_{\pm 0.43}^{*}$ 38.60 $43.38_{\pm 0.47}^{*}$ $33.56_{\pm 0.44}^{*}$ 37.84
Qwen2.5-7B $55.77_{\pm 0.43}^{*}$ $39.45_{\pm 0.45}^{*}$ 46.21 $45.36_{\pm 0.47}^{*}$ $49.82_{\pm 0.47}^{*}$ 47.49
Qwen2.5-72B $62.58_{\pm 0.44}$ $43.66_{\pm 0.44}$ 51.43 $74.95_{\pm 0.40}$ $51.45_{\pm 0.46}$ 61.02
SFT Llama-3.2-1B $65.29_{\pm 0.39}$ $52.92_{\pm 0.47}$ 58.46 $71.58_{\pm 0.39}$ $62.24_{\pm 0.46}$ 66.59
Llama-3.2-3B $67.00_{\pm 0.39}$ $56.57_{\pm 0.45}$ 61.35 $75.72_{\pm 0.37}$ $64.64_{\pm 0.45}$ 69.74
Qwen2.5-1.5B $63.81_{\pm 0.40}$ $57.25_{\pm 0.45}$ 60.35 $76.35_{\pm 0.37}$ $69.85_{\pm 0.44}$ 72.96
Qwen2.5-3B $68.77_{\pm 0.38}$ $58.42_{\pm 0.45}$ 63.18 $76.89_{\pm 0.37}$ $\textbf{70.74}_{\pm 0.43}$ 73.69
Qwen2.5-7B $\textbf{70.75}_{\pm 0.39}$ $\textbf{60.31}_{\pm 0.44}$ 65.11 $\textbf{80.15}_{\pm 0.34}$ $69.20_{\pm 0.44}$ 74.27
  • *” denotes the result is significantly worse than results of corresponding SFT model in t-test with $p<0.05$.
5.1 Setups
Dataset Splitting. We split the English data at an 8:2 ratio to obtain the training and test set and follow the dataset splits from the original data sources for Chinese data. The detailed training and test set statistics are shown in Appendix E. During the training process, we balance the number of samples with and without privacy leakage in the training data, ensuring a ratio of approximately 1:1.
Models. We take the instruction-tuned version of different series of language models with various scales, including Llama-3.2-1B and 3B llama3 , and Qwen-2.5-1.5B, 3B, 7B, and 72B qwen2 . All the pretrained checkpoints are loaded from huggingface [^2] https://huggingface.co.
Implementation Details. We implement the Privacy Detection Model and all the baselines in PyTorch. The workstation has 8 NVIDIA V100 32G GPUs. To improve training efficiency, we adopt LoRA lora for tuning our model. More training settings are listed in Appendix H.
Figure 6: Performance of different privacy detection baselines with information-level metrics (%).
5.2 Local Privacy Detection Baselines
Tuning-free Baselines Tuning-free methods leverage the inherent capabilities of pre-trained models for privacy detection. (1) Zero-shot Generation (ZG) gpt3 . We directly prompt local LLMs to determine whether the query leaks privacy or to extract privacy phrase and corresponding information. In the prompt, we explicitly specify that privacy information in user dialogue includes various aspects such as opinions, preferences, intentions, etc. We also require that the recognized privacy be relevant to the user and ask the model to think step by step. Finally, we guide the model to output structured results in the specified format. (2) In-Context Learning (ICL) icl . Based on the prompts of ZG, we further insert five examples from training data based on the template for inference with local privacy detection LLM. After the examples, the current user query is added for privacy detection. All corresponding prompts are provided in Appendix G.
Tuning-based Baseline. We use Supervised Fine-Tuning (SFT) with instructions to enable the model to judge privacy leakage or to generate structured phrases and corresponding information based on the given user’s conversational content. Additionally, we find that overly long instruction templates will reduce the model’s sensitivity to user queries, resulting in a performance drop. The instruction template used during training is provided in Appendix G.
5.3 Main Results
Privacy Leakage Classification. We first evaluate the performance of different baselines on the binary classification task and the results are presented in Table 3. We can find that model performance generally improves with an increase in parameter size. Notably, the largest model in our evaluation, Qwen2.5-72B, excels with ZG, achieving 14.08% and 18.24% improvement for query-level accuracy compared to Qwen2.5-1.5B in English and Chinese dataset, respectively. By comparing SFT with ZG and ICL across different models, we can find that the fine-tuned Qwen2.5-7B achieves the best performance on both English and Chinese datasets. Specifically, SFT with Qwen2.5-7B outperforms ZG and ICL with the same model by over 14.59% and 15.35%, respectively.
Privacy Phrase Extraction. We then investigate the quality of extracted privacy phrases by different baselines. From Table 3, we can also observe that ICL outperforms ZG in most cases. For example, with the Qwen2.5-7B, ICL achieves 3.66% and 3.82% improvement for phrase-level F1 over ZG with the same model in English and Chinese dataset, respectively. These results suggest that inserting examples from the training set enhances privacy detection. Additionally, by fine-tuning on our dataset, the 1B model outperforms the direct inference results of the 72B model by 7.03% and 5.57% in English and Chinese dataset respectively, which is friendly for local devices. Furthermore, the fine-tuned Qwen2.5-7B achieves an improvement of 13.68% and 13.25% in phrase-level F1 compared to Qwen2.5-72B with ICL in English and Chinese dataset, respectively. These significant gains underscore the importance of tuning on our custom dataset.
Privacy Information Summarization. Figure 6 presents the performance of different local privacy detection baselines for information-level F1. We can find that by supervised fine-tuning with the training set, Qwen2.5-7B achieves the best performance on different datasets. Specifically, the fine-tuned Qwen2.5-7B outperforms ZG with Qwen2.5-7B by 21.06% and 28.03%, and outperforms ICL with Qwen2.5-7B by 16.80% and 20.33% in English and Chinese dataset, respectively. We can also observe that fine-tuned 1B model outperforms both ZG and ICL with 72B model by over 10.83% and 7.18% in English and Chinese dataset, respectively. However, there still a gap between the performance of more light-weight LLMs and the ground truth in fine-grained privacy extraction. This emphasizes the need for exploring more effective methods grounded in our dataset. More detailed discussion is provided in Appendix I. Furthermore, to explore the real-time inference efficiency of the local fine-tuned models, we measured the average privacy detection time per query in Appendix J.
5.4 Case Study
Table 4: Case study with Qwen2.5-7B-Instruct.
Query :( I am worried I will screw up the computer because i am bad on coding.
ZG Phrase 1: I am bad on coding
Info 1: The user is worried about their coding abilities.
ICL Phrase 1: bad on coding
Info 1: The user expresses concern about their coding abilities and fears making mistakes that could damage a computer.
SFT Phrase 1: I will screw up the computer
Info 1: The user worries about making mistakes with the computer.
Phrase 2: I am bad on coding
Info 2: The user perceives themselves as not skilled in coding.
To better assess the performance of various local privacy detection methods, we present a case study of phrase and information extraction in Table 4. We can observe that both ZG and ICL fail to capture the user’s specific concerns, while ICL generates more detailed privacy information compared to ZG. In contrast, SFT successfully identifies all relevant privacy phrases from the user input and generates accurate privacy information. We also present a bad case for further analysis of these baselines’ performance in Appendix K.
6 Conclusion, Impact and Future Work
In this work, we have studied privacy detection for real-name LLM interaction. We have constructed the first large-scale multilingual dialogue dataset for different privacy detection tasks, achieved through an automated scalable privacy annotation pipeline for privacy phrases extraction and corresponding information annotation. Additionally, we have developed evaluation metrics at query-level, phrase-level, and information-level. Finally, we have evaluated various local privacy detection methods, revealing that privacy detection in LLM interactions remains a challenging task. As for future research opportunity, the dataset provides a valuable foundation to explore more effective privacy detection methods that can be efficiently deployed on different local devices.
Limitations and Discussion. We acknowledge that the built dataset is designed for privacy detection of user-related information in user’s daily interactions with LLM and should not be regarded as a comprehensive dataset, as it may not suitable for all possible privacy leakage scenarios. For example, detecting privacy from user’s query whether it leaks any information from enterprise database. Extending the dataset by applying the built automated annotation pipeline to different private scenarios remains a future prospect. Ethical statement is provided in Appendix M.
Appendix A Prompts for Automated Privacy Annotation Pipeline
Table 10, Table 11 - 12, Table 13 - 16, and Table 17 provides prompts for privacy leakage classification, extensive privacy categories extraction, privacy phrase extraction, and privacy information annotation in designed pipeline, respectively.
Appendix B Introduction of Data Source
ShareGPT sharegpt is composed of dialogues between users and ChatGPT, with 79K raw dialogues and 243K user queries. CrossWOZ crosswoz is a task-oriented dialogue dataset covering five domains, containing 6K dialogues and 47K user queries. DuConv duconv focuses on movie and film star topics, with 22K dialogues and 99K user queries. Both CrossWOZ and DuConv data are collected through dialogues conducted by two workers, based on specific topics or knowledge backgrounds. The LCCC-base lccc is derived from the Weibo corpus. We selected a subset of this dataset as our source data, including 50K dialogues and 76K user queries.
Appendix C Dataset Examples
Table 18 shows raw data examples and examples of the constructed dataset. One data sample is consists of a user-initiated query and a response from the assistant. Based on this, we add privacy detection results to each query, in the form of a list of privacy. Each element in the list consists of a phrase extracted from the query and the corresponding specific privacy information. In terms of dataset processing, we also retain the contextual relationship between the original queries in one dialogue.
Table 5: The comparison with existing dataset.
Dataset Target Labeling Data Size Information Language
I2B2/UTHealth i2b2 Personal Health Manual 1,304 $\times$ EN
CodE ALLTAG codealltag Email Manual 2,309 $\times$ DE
TAB tab Court Judgments Manual 1,268 $\times$ EN
Reddit-SD selfdisclosures Comments Manual 2,415 $\times$ EN
Ours LLM Interactions Automated 249,683 $\surd$ EN, ZH
Appendix D Comparison with Existing Datasets
As shown in Table 5, our dataset is centered on user interactions, where the definition of privacy extends beyond identifiable details and encompasses a broader range of sensitive information. Additionally, our dataset is annotated automatically, resulting in significantly larger scale compared to existing datasets. Furthermore, it offers detailed information for each extracted privacy phrase.
Appendix E Details on Training and Testing Sets
Table 6: Statistics of the dataset with automatically annotated privacy phrases and information.
Language Training set Test set
# Non-Leak # Leak # Phrase # Non-Leak # Leak # Phrase
English 52,053 26,156 67,780 12,792 6,658 17,540
Chinese 98,100 38,025 60,611 10,633 5,230 8,299
Table 6 lists detailed statistics of the dialogue dataset with annotated privacy phrases and information. The privacy phrase set and information set is empty for samples without privacy leakage, while the average length of phrase set and information set for samples with privacy leakage is 2.6 in English part and 1.6 in Chinese part.
Appendix F Human Annotation Guidelines
F.1 Annotation Instructions for Privacy Leakage
Annotators are eight students in the lab. They are instructed to determine whether the query reveals privacy information about the user who posed it. Given that the user’s identifier is exposed, annotators must assess whether the query discloses any new privacy information about the user or related individuals or entities. Privacy information includes opinions, concerns, preferences, activities, intentions, or any other information that could be considered private or sensitive based on the query’s context. Annotators’ responses should only be “true” or “false.”
F.2 Annotation Instructions for Privacy Phrase Extraction
Based on the definition of privacy information in user dialogues, annotators are tasked with extracting phrases that reveal private details from the query. These extracted phrases should be directly copied from the query without any alterations to their format or wording. The selected phrases must clearly identify and reflect the user’s private information, serving as key indicators of such information. Each annotated result is reviewed for accuracy by a second annotator to ensure consistency and reliability.
Appendix G Prompts for Local Privacy Detection Methods
Table 19 - Table 21, Table 22 - 24, and Table 25 - 27 shows detailed prompts used for zero-shot generation, in-context learning, and training privacy detection models, respectively. For zero-shot generation and in-context learning, we prompt language models to extract privacy phrases from user dialogues and generate specific information for each identified phrase. For fine-tuning, we employ a concise template to guide the model, allowing it to learn both the content and output format from the training data.
Appendix H Additional Training Details
We utilize the AdamW optimization scheme with PyTorch for all methods. For Llama3.2-1B-Instruct and Qwen2.5-1.5B-Instruct, the learning rate is set between a maximum of 2e-3 and a minimum of 1e-3. For models larger than 1.5B parameters, the learning rate is set to a maximum of 2e-4 and a minimum of 1e-4. Training is conducted with a batch size of 4, for 10 epochs, and a maximum token length of 1024. For LoRA, we configure the rank to 16, apply a dropout rate of 0.05, and perform weight decomposition on the query, key, and value matrices, as well as the dense MLP layers in the transformer blocks. Additionally, during training, the parameters of Qwen and Llama are set to FP32 to maintain precision.
Appendix I Discussion of Main Results
The main purpose of local privacy detection methods is to perform privacy leakage classification, phrase extraction, or information summarization without the need to upload raw user queries online. Although the fine-tuned Qwen2.5-7B model achieves an accuracy of 87.56% for binary privacy leakage classification task, its performance on more complex tasks can not meet the requirements for practical implementation. For example, the phrase-level F1 score for the best fine-tuned model is 74.27% and the best information-level F1 score is 44.70% on the Chinese dataset. This also highlights a gap between the performance of more light-weight LLMs and the ground truth in fine-grained privacy extraction. To address this, a research line is to explore new methods to enhance the performance of generative locally deployable LLMs on these tasks, including the design of novel model architectures and the incorporation of different training methods, such as reinforcement learning.
Appendix J Inference Time for Local Privacy Detection Models
Table 7: Average privacy detection time per query (s) of fine-tuned local models for English samples.
Local Device Llama3.2-1B Qwen2.5-1.5B Qwen2.5-3B Llama3.2-3B Qwen2.5-7B
GeForce RTX 2080 Ti 0.1772 0.2113 0.2668 0.4062 out-of-memory
GeForce RTX 4090 0.1191 0.1209 0.1630 0.2025 0.3980
Table 8: Average privacy detection time per query (s) of fine-tuned local models for Chinese samples.
Local Device Llama3.2-1B Qwen2.5-1.5B Qwen2.5-3B Llama3.2-3B Qwen2.5-7B
GeForce RTX 2080 Ti 0.0885 0.0922 0.1279 0.1623 out-of-memory
GeForce RTX 4090 0.0521 0.0623 0.0789 0.0879 0.1754
To explore the real-time inference efficiency of the local privacy detection models which are fine-tuned on our dataset, we measured the average privacy detection time per query for both English and Chinese samples. The model precision is set to FP16 and the results is shown in Table 7 and Table 8. These results reveal the time efficiency of local privacy detection with billion-scale LLMs.
Appendix K Case Study
Table 9: A bad case of Qwen2.5-7B-Instruct with SFT.
Query My 3 year old repeats syllables twice, rather than pronouncing both syllables of the word. For example he says ‘The taitai needs wawa’ instead of ‘the tiger needs water.’ He sounds a bit babyish. Why does he do this? How can I help him? Should I be worried about it?
SFT Phrase 1: repeats syllables twice
Information 1: The user is concerned about their 3-year-old child’s speech development, specifically regarding the pronunciation of syllables.
Phrase 2: Should I be worried about it?
Information 2: The user is concerned about their 3-year-old child’s speech development.
GT Phrase 1: My 3 year old
information 1: The user’s child is 3 years old.
Phrase 2: He sounds a bit babyish.
Information 2: The user’s child is described as having speech that sounds babyish, which suggests a concern about speech development.
Phrase 3: repeats syllables twice
Information 3: The user’s 3-year-old child has a speech pattern of repeating syllables twice rather than pronouncing both syllables of words.
Current local privacy detection methods also have some limitations. Table 9 presents a bad case from the fine-tuned Qwen2.5-7B, where the local model incorrectly extracted a general question that did not actually expose the user’s privacy. Moreover, the model failed to identify certain implicit privacy phrases. These shortcomings highlight the challenges of extracting phrases and generating accurate information. Light-weight LLMs, with limited text comprehension abilities, require more sophisticated methods to be developed in future work for effective privacy detection on the user side.
Appendix L API Cost
We use gpt-4o-2024-08-06 as the base LLM for our automated privacy annotation pipeline, which extracts privacy phrases and annotates information. For the English dataset, the total token count is 296M tokens for input and 152M for output. For the Chinese dataset, the total token count is 219M tokens for input and 99M for output. The total annotation cost amounts to approximately \$3,800 and the average cost is \$0.015 per sample.
Appendix M Ethical Statement
We investigates privacy detection in user’s real-name interactions with LLM. The design, implementation, and use of the automated privacy annotation pipeline, along with the local privacy detection methods, are guided by ethical principles for responsible AI deployment. (1) Dataset Construction: All dialogue data used in this study are publicly available and properly cited, with no ethical concerns regarding their use for research purposes. (2) Fair Usage: The primary goal of the privacy annotation pipeline is to extract privacy phrases and annotate privacy information from public dialogue datasets, rather than conducting malicious analysis of user privacy. We believe that the responsible application of our design aligns with ethical guidelines.
Table 10: Prompt template for privacy leakage classification for English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for privacy leakage classification (English).
[…]
A user poses a query
"<|INPUT|>"
to you.
Before the user poses the query, you only know the real unique identifier of the user, and do not know any other privacy information of the user.
After receiving the query with the user’s real unique identifier, think step by step to judge whether the query reveals any new privacy information of the user or the user’s related people or things to you.
The privacy information should be relevant to the user or the user’s related people or things based on the context of the query.
The privacy information includes any opinions, concerns, preferences, activities, intentions, or any other information that could be considered private or sensitive based on the context of the query.
Give your think steps. Your reasoning must depend on the direct evidences contained in the query and cannot be based on assumption.
Then, output a json as follows:
“‘json
    "reason": reason for your judgment,
    "judgment": true or false
Here are some examples: User’s query: We want to play football today. JSON Output: “‘json
    "reason": "This query exposes the activity of the user and people related to the user intend to do.",
    "judgment": true
…… (other examples)
Prompts for privacy leakage classification (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
{CJK}UTF8gbsn "<|INPUT|>"
{CJK}UTF8gbsn
{CJK}UTF8gbsn 在用户提出请求之前,你只知道用户的真实的唯一标识符,并且不知道用户的任何其他的隐私敏感信息。
{CJK}UTF8gbsn 在收到带有用户真实的唯一标识符的请求后,请你进行逐步的思考(Think Step By Step)来判断该请求是否向你透露了任何新的关于用户或与用户相关的人或事的隐私敏感信息
{CJK}UTF8gbsn
{CJK}UTF8gbsn 隐私敏感信息应与用户或与用户相关的人或事相关,该相关性体现在用户请求的上下文中。
{CJK}UTF8gbsn 隐私敏感信息包括任何观点、担忧、偏好、活动、意图或任何其他根据查询上下文可被视为私密或敏感的信息。
{CJK}UTF8gbsn
{CJK}UTF8gbsn 给出你的思考与分析步骤。你的推理必须依赖于查询中包含的直接证据,不能基于假设
{CJK}UTF8gbsn
{CJK}UTF8gbsn 最后,输出如下格式的json:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "reason": 你判断的理由,
{CJK}UTF8gbsn     "judgment": true 或 false
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn
{CJK}UTF8gbsn 以下是一些示例:
{CJK}UTF8gbsn 用户的请求:我们今天想踢足球。
{CJK}UTF8gbsn JSON输出:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "reason": "该请求透露了用户及与用户相关的人打算进行的活动。",
{CJK}UTF8gbsn     "judgment": true
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn …… (other examples)
Table 11: Prompt template for privacy categories extraction in step 2 for English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for privacy categories extraction (English).
[…]
Given a user’s query
"<|INPUT|>"
Identify which phrases in the query leak the user’s privacy information and the correponding categories of privacy information.
Let’s think step by step and output a json finally.
The result output should be in json format with the following format:
“‘json
    privacy information category 1: [phrase 1, phrase 3, …],
    privacy information category 2: [phrase 2, phrase 4, …]
Here is an example
"Query": "I’m a teacher at China, majoring in computer science. I will go back home to London."
Output:
“‘json
    "Job": ["teacher"],
    "Subject": ["computer science"],
    "Location": ["China", "London"]
Prompts for privacy categories extraction (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
{CJK}UTF8gbsn "<|INPUT|>"
{CJK}UTF8gbsn
{CJK}UTF8gbsn 提取请求中泄露用户隐私信息的短语,并将其分类为相应的隐私信息类别
{CJK}UTF8gbsn 让我们一步一步地思考,最后输出一个json,格式如下:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "隐私信息类别1": ["短语1", "短语3", …],
{CJK}UTF8gbsn     "隐私信息类别2": ["短语2", "短语4", …]
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn
{CJK}UTF8gbsn 以下是一个示例:
{CJK}UTF8gbsn 用户的请求:"我是中国的一名教师,主修计算机科学。我将回家到伦敦。"
{CJK}UTF8gbsn JSON输出:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "职业": ["教师"],
{CJK}UTF8gbsn     "专业": ["计算机科学"],
{CJK}UTF8gbsn     "地点": ["中国", "伦敦"]
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
Table 12: Prompt template for categories deduplication in step 2 for English and Chinese dataset, where <|INPUT|> is the input of privacy categories and corresponding phrase examples, and <|BLOCK-SIZE|> is the number of input categories.
Prompts for categories deduplication (English).
[…]
Given the following <|BLOCK-SIZE|> privacy categories and the list of example phrases belong to each privacy category:
<|INPUT|>
If privacy category A and privacy category B have the same meaning, just keep category A.
Let’s think step by step and output a json list finally.
Output the new privacy categories after deduplication in the format of a list in json, for example:
[privacy category 1, privacy category 2, … ]
Prompts for categories deduplication (Chinese).
[…]
{CJK}UTF8gbsn 给定以下<BLOCK-SIZE>个隐私类别及每个隐私类别对应的示例短语列表:
{CJK}UTF8gbsn <|INPUT|>
{CJK}UTF8gbsn
{CJK}UTF8gbsn 对隐私类别进行去重,如果隐私类别A和隐私类别B具有相同的含义,只保留类别A
{CJK}UTF8gbsn 让我们一步一步地思考,最后输出一个json列表。
{CJK}UTF8gbsn 以列表格式输出去重后的新的隐私类别,例如:
{CJK}UTF8gbsn [隐私类别1, 隐私类别2, … ]
Table 13: Prompt template for privacy phrase extraction with categories in step 3 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|CATEGORIES|> is the input of privacy categories.
Prompts for privacy phrase extraction (English).
[…]
A user poses a query
"<|INPUT|>"
to you.
Extract informative phrases from the query that definitely reveal the privacy-sensitive information of the user, or user related people or things.
The phrases must be included exactly as they appear in the query’s content. Do not alter the format or wording of the phrases. For example, spelling errors, number of spaces, capitalization, etc., must all be preserved exactly same as the query’s content.
Give your think steps. The reasoning must depend on the direct evidences contained in the query and cannot be based on assumption.
Then, match each extracted phase with the most appropriate privacy category in the following pre-defined list, and also note if the extracted phrase is not matched with any given privacy category, it should be excluded.
Privacy Categories: <|CATEGORIES|>
The result output should be in json format with the following format:
    "phrase 1" : "privacy category 3",
    "phrase 2" : "privacy category 4"
Here are some examples
User’s Query: "I’m a teacher at China, majoring in computer science. I will go back home to London."
Privacy Categories: ["Name", "Age", "Gender", "Job", "Subject", "Location"]
JSON Output:
“‘json
    "teacher": "Job",
    "computer science": "Subject",
    "China": "Location",
    "London": "Location"
…… (other examples)
Prompts for privacy phrase extraction (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
{CJK}UTF8gbsn "<|INPUT|>"
{CJK}UTF8gbsn 从请求的内容中提取短语,该短语明确揭示用户的,或与用户相关的人或事物的隐私敏感信息。
{CJK}UTF8gbsn 短语必须与请求内容中出现的完全一致。不要更改短语的格式或措辞,即拼写错误、空格数量、大小写等,都必须与请求内容完全相同。
{CJK}UTF8gbsn 给出你的思考步骤(Think Step by Step)。推理必须依赖于请求中包含的直接证据,不能基于假设
{CJK}UTF8gbsn 然后,将每个提取的短语与以下预定义列表中最合适的隐私类别匹配,并且如果提取的短语未与任何给定的隐私类别匹配,则应将其排除。
{CJK}UTF8gbsn
{CJK}UTF8gbsn 隐私类别:<|CATEGORIES|>
{CJK}UTF8gbsn
{CJK}UTF8gbsn 结果输出应为json格式,格式如下:
[…]
{CJK}UTF8gbsn     "短语1" : "隐私类别3",
{CJK}UTF8gbsn     "短语2" : "隐私类别4"
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn
{CJK}UTF8gbsn 以下是一些示例
{CJK}UTF8gbsn 用户的请求:"我是中国的一名教师,主修计算机科学。我将回家到伦敦。"
{CJK}UTF8gbsn 隐私类别:["姓名", "年龄", "性别", "职业", "专业", "地点"]
{CJK}UTF8gbsn JSON输出:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "教师": "职业",
{CJK}UTF8gbsn     "计算机科学": "专业",
{CJK}UTF8gbsn     "中国": "地点",
{CJK}UTF8gbsn     "伦敦": "地点"
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn …… (other examples)
Table 14: Prompt template for privacy phrase deduplication in step 3 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|PHRASES|> is the input of extracted privacy phrase.
Prompts for privacy phrase deduplication (English).
[…]
Given a user’s query and privacy phrases extracted from the query,
Query: "<|INPUT|>"
Privacy Phrases: <|PHRASES|>
Think step by step to deduplicate the privacy phrases by strictly following the given rules:
[Rule 1] If privacy phrase A and privacy phrase B have the same meaning, keep phrase which is more concise and clear.
[Rule 2] If privacy phrase A is part of privacy phrase B, keep phrase which has more information and is clear.
Give your think steps. Then output the new privacy phrases after deduplication in the format of a list in json, for example:
“‘json
[new privacy phrase 1, new privacy phrase 2, ]
Here is a examples:
Query: "We are happy and plan to play football."
Privacy Phrases: ["happy", "We are happy", "plan to play football", "football", "play football"]
JSON output:
“‘json
["happy", "play football"]
Prompts for privacy phrase deduplication (Chinese).
[…]
{CJK}UTF8gbsn 给定一个用户请求和从请求中提取的隐私短语,
{CJK}UTF8gbsn 用户请求:"<|INPUT|> "
{CJK}UTF8gbsn 隐私短语:<|PHRASES|>
{CJK}UTF8gbsn
{CJK}UTF8gbsn 一步一步地思考(Think Step by Step),严格按照给定的规则对隐私短语进行去重
{CJK}UTF8gbsn [规则1] 如果隐私短语A和隐私短语B具有相同的含义,保留更简洁明了的短语。
{CJK}UTF8gbsn [规则2] 如果隐私短语A是隐私短语B的一部分,保留信息量更大且清晰的短语。
{CJK}UTF8gbsn 给出你的思考步骤。然后以json列表格式输出去重后的新隐私短语,例如:
{CJK}UTF8gbsn “‘json
{CJK}UTF8gbsn [新隐私短语1, 新隐私短语2, ]
Table 15: Prompt template for filtering privacy phrase with rule 1 in step 3 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|PHRASE|> is the input of the phrase.
Prompts for filtering privacy phrase with rule 1 (English).
[…]
A user poses a query
"<|INPUT|>"
to you with a real unique identifier.
Given a phrase extracted from the query as follows:
Phrase: "<|PHRASE|>"
Think step by step to reason about whether the phrase satisfies the following rule:
[Rule] the phrase directly links to the user or user’s related people or things in the context of the user’s query.
Give your think steps. The reasoning must depend on the direct evidences contained in the query and cannot be based on assumption.
Then, output a json as follows:
“‘json
    "reason" : reason for your judgment,
    "judgment" : true or false
Here are some examples:
User’s Query: "We plan to play football with Ross."
Phrase: "play football with Ross"
JSON Output:
“‘json
    "reason" : "Playing football is the user’s plan that related to the user. Ross is also related to the user.",
    "judgment" : true
…… (other examples)
Prompts for filtering privacy phrase with rule 1 (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
{CJK}UTF8gbsn "<|INPUT|>"
{CJK}UTF8gbsn 并给定用户的真实的身份标识
{CJK}UTF8gbsn 给定从请求中提取出短语如下:
{CJK}UTF8gbsn 短语:" <|PHRASE|> "
{CJK}UTF8gbsn
{CJK}UTF8gbsn 一步一步地思考(Think Step by Step),判断该短语是否满足以下规则
{CJK}UTF8gbsn [规则] 该短语在用户请求的上下文中直接与用户或与用户相关的人或事物相关联。
{CJK}UTF8gbsn 给出你的思考步骤。推理必须依赖于请求中包含的直接证据,不能基于假设
{CJK}UTF8gbsn 然后,输出如下格式的json:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "reason" : 你判断的理由,
{CJK}UTF8gbsn     "judgment" : true 或 false
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn
{CJK}UTF8gbsn 以下是一些示例:
{CJK}UTF8gbsn 用户的请求:"我们计划和罗斯一起踢足球。"
{CJK}UTF8gbsn 短语:"和罗斯一起踢足球"
{CJK}UTF8gbsn JSON输出:
{CJK}UTF8gbsn “‘json
[…]
{CJK}UTF8gbsn     "reason" : "踢足球是用户的计划,与用户相关。罗斯也与用户相关。",
{CJK}UTF8gbsn     "judgment" : true
{CJK}UTF8gbsn }}\}}
{CJK}UTF8gbsn ”’
{CJK}UTF8gbsn …… (other examples)
Table 16: Prompt template for filtering privacy phrase with rule 2 in step 3 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|PHRASE|> is the input of the phrase.
Prompts for filtering privacy phrase with rule 2 (English).
[…]
A user poses a query
"<|INPUT|>"
to you with a real unique identifier.
Given a phrase extracted from the query as follows:
Phrase: "<|PHRASE|>"
Think step by step to reason about whether the phrase satisfies the following rule:
[Rule] the phrase should not be a general phrase that provide NO information related to the user based on the context of the user’s query.
Give your think steps. The reasoning must depend on the direct evidences contained in the query and cannot be based on assumption.
Then, output a json as follows:
“‘json
    "reason" : reason for your judgment,
    "judgment" : true or false
Here are some examples:
User’s Query: "I want to go a place."
Phrases: "place"
JSON Output:
“‘json
    "reason" : "The ‘place’ term has no clear reference.",
    "judgment" : false
…… (other examples)
Prompts for filtering privacy phrase with rule 2 (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
"<|INPUT|>"
{CJK}UTF8gbsn 并给定用户的真实的身份标识
{CJK}UTF8gbsn 给定从请求中提取出短语如下:
{CJK}UTF8gbsn 短语:" <|PHRASE|> "
{CJK}UTF8gbsn 一步一步地思考(Think Step by Step),判断该短语是否满足以下规则
{CJK}UTF8gbsn [规则] 该短语不应是一个在用户请求上下文中不提供与用户相关信息的一般短语。
{CJK}UTF8gbsn 给出你的思考步骤。推理必须依赖于查询中包含的直接证据,不能基于假设
{CJK}UTF8gbsn 然后,输出如下格式的json:
“‘json
    "reason" : {CJK}UTF8gbsn 你判断的理由 ,
    "judgment" : true {CJK}UTF8gbsn 或 false
{CJK}UTF8gbsn 以下是一些示例:
{CJK}UTF8gbsn 用户的请求:"我想去一个地方。"
{CJK}UTF8gbsn 短语:"地方"
{CJK}UTF8gbsn JSON输出:
“‘json
{CJK}UTF8gbsn     "reason" : "‘地方’一词没有明确的指代。" ,
    "judgment" : false
…… (other examples)
Table 17: Prompt template for privacy information annotation in step 4 for English and Chinese dataset, where <|INPUT|> is the input of user query and <|PHRASE|> is the input of the phrase.
Prompts for privacy information annotation (English).
[…]
A user poses a query
"<|INPUT|>"
to you with a real unique identifier.
Given a phrase extracted from the query as follows:
Phrase: "<|PHRASE|>"
The phrase reveals new privacy-sensitive information of the user, or user related people or things to you.
Think step by step.
Summarize the specific privacy information the phrase has leaked.
Then, output a json as follows:
    "phrase": PHRASE,
    "privacy information": PRIVACY-INFORMATION
Here are some examples:
User’s Query: "I am preparing to go back home to Minhang District, Shanghai."
Phrase: "Minhang District, Shanghai"
JSON Output:
“‘json
    "phrase": "No. 800, Dongchuan Road, Minhang District, Shanghai",
    "privacy information": "The user’s home address is No. 800, Dongchuan Road, Minhang District, Shanghai."
…… (other examples)
Prompts for privacy information annotation (Chinese).
[…]
{CJK}UTF8gbsn 用户向你提出一个请求
"<|INPUT|>"
{CJK}UTF8gbsn 并给定用户的真实的身份标识
{CJK}UTF8gbsn 给定从请求中提取出短语如下:
{CJK}UTF8gbsn 短语: "<|PHRASE|>"
{CJK}UTF8gbsn 该短语向你揭示了用户或与用户相关的人或事物的,新的隐私敏感信息
{CJK}UTF8gbsn 一步一步地思考(Think Step by Step)。
{CJK}UTF8gbsn 并总结短语泄露的具体隐私敏感信息
{CJK}UTF8gbsn 然后,输出如下格式的JSON:
“‘json
    "phrase": {CJK}UTF8gbsn 短语 ,
    "privacy information": {CJK}UTF8gbsn 隐私信息
{CJK}UTF8gbsn 以下是一些示例:
{CJK}UTF8gbsn 用户的请求:"我正在准备回家,地址是上海市闵行区东川路800号。"
{CJK}UTF8gbsn 短语:"上海市闵行区东川路800号"
{CJK}UTF8gbsn JSON输出:
“‘json
    "phrase": "{CJK}UTF8gbsn 上海市闵行区东川路800号" ,
    "privacy information": "{CJK}UTF8gbsn 用户的家庭地址是上海市闵行区东川路800号。"
…… (other examples)
Table 18: Data samples with privacy phrase and privacy information. The samples are come from one raw dialogue.
Raw data examples
[…]
"id": "ShareGPT-pKNQqpRRE1",
"conversation": [
        "user": "Can you please list the operators know to work for ArcelorMittal in the world and who are not OBS?",
        "assistant": "As an AI language model, I don’t have access to up-to-date information ……",
        "user": "David Glijer is a key target for us. How would you recommand to engage him?",
        "assistant": "When engaging with David Glijer, it’s important to understand ……",
Annotated data examples of our dataset
[…]
"id": "ShareGPT-pKNQqpRRE1",
"conversation": [
        "user": "Can you please list the operators know to work for ArcelorMittal in the world and who are not OBS?",
        "assistant": "As an AI language model, I don’t have access to up-to-date information ……",
        "privacy": []
        "user": "David Glijer is a key target for us. How would you recommand to engage him?",
        "assistant": "When engaging with David Glijer, it’s important to understand ……",
        "privacy": [
                "phrase": "David Glijer",
                "privacy information": "David Glijer is identified as a key target of interest, suggesting a strategic focus in a business context."
                "phrase": "key target for us",
                "privacy information": "David Glijer is an important person of interest or focus for the user’s objectives, implying a strategic or professional engagement."
                "phrase": "engage him",
                "privacy information": "The user has plans to strategically engage or interact with David Glijer, indicating he is a subject of interest."
Table 19: Prompt template for zero-shot generation of privacy leakage classification task in English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for zero-shot generation (English).
[…]
Determine whether the query leaks the privacy information of the user, or user related people or things.
The privacy information includes any opinions, concerns, preferences, activities, intentions, or any other information that could
be considered private or sensitive based on the context of the query.
Finally, output your answer in JSON.
Here is the format:
“‘json
    "judgment": true or false
The user’s QUERY is
"<|QUERY|>"
Prompts for zero-shot generation (Chinese).
[…]
{CJK}UTF8gbsn 判断用户请求是否泄露了用户、用户相关人员或事物的隐私敏感信息。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。
{CJK}UTF8gbsn 以下是格式:
“‘json
    "judgment": true or false
{CJK}UTF8gbsn 用户的请求是:
"<|QUERY|>"
Table 20: Prompt template for zero-shot generation of privacy phrase extraction task in English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for zero-shot generation (English).
[…]
Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
The privacy information includes any opinions, concerns, preferences, activities, intentions, or any other information that could be considered private or sensitive based on the context of the query.
The phrase should be relevant to the user or the user’s related people or things based on the context of the query.
The phrase should not be a general term.
You should think step-by-step.
Finally, output your answer in JSON. If there is no phrase extracted, the output JSON list should be empty.
Here is the format:
“‘json
["phrase1", "phrase2", …]
The user’s QUERY is
"<|QUERY|>"
Prompts for zero-shot generation (Chinese).
[…]
{CJK}UTF8gbsn 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性
{CJK}UTF8gbsn 隐私短语应当有具体的指代,而不是一个泛指的词汇。
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
{CJK}UTF8gbsn 以下是格式:
“‘json
{CJK}UTF8gbsn ["短语1", "短语2", …]
{CJK}UTF8gbsn 用户的请求是:
"<|QUERY|>"
Table 21: Prompt template for zero-shot generation of privacy information summarization task in English and Chinese dataset, where <|INPUT|> is the input of user query.
Prompts for zero-shot generation (English).
[…]
Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
The privacy information includes any opinions, concerns, preferences, activities, intentions, or any other information that could
be considered private or sensitive based on the context of the query.
The phrase should be relevant to the user or the user’s related people or things based on the context of the query.
Then for each extracted phrase, summarize the specific privacy information the phrase has leaked.
You should think step-by-step.
Finally, output your answer in JSON. If there is no phrase extracted, the output JSON list should be empty.
Here is the format:
“‘json
        "phrase": PHRASE-1,
        "privacy information": PRIVACY-INFORMATION-1
The user’s QUERY is
"<|QUERY|>"
Prompts for zero-shot generation (Chinese).
[…]
{CJK}UTF8gbsn 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
{CJK}UTF8gbsn 然后,对于每个提取的短语,总结该短语泄露的具体隐私信息
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
{CJK}UTF8gbsn 以下是格式:
“‘json
{CJK}UTF8gbsn         "phrase": "短语",
{CJK}UTF8gbsn         "privacy information": "具体的隐私信息"
{CJK}UTF8gbsn 用户的请求是:
"<|QUERY|>"
Table 22: Prompt template for in-context learning of privacy leakage classification task in English and Chinese dataset, where <|INPUT|> is the input of user query and <|CASE|>is structured examples from training set.
Prompts for in-context learning (English).
[…]
Determine whether the query leaks the privacy information of the user, or user related people or things.
The privacy information includes any opinions, concerns, preferences, activities, intentions, or any other information that could
be considered private or sensitive based on the context of the query.
Finally, output your answer in JSON.
Here are some examples:
<|CASE|>End of examples.
The user’s current Query is
"<|QUERY|>"
Prompts for in-context learning (Chinese).
[…]
{CJK}UTF8gbsn 判断用户请求是否泄露了用户、用户相关人员或事物的隐私敏感信息。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。
{CJK}UTF8gbsn 以下是一些例子:
{CJK}UTF8gbsn <|CASE|>例子结束。
{CJK}UTF8gbsn 用户当前的请求是:
"<|QUERY|>"
Table 23: Prompt template for in-context learning of privacy phrase extraction task in English and Chinese dataset, where <|INPUT|> is the input of user query and <|CASE|>is structured examples from training set.
Prompts for in-context learning (English).
[…]
Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
The privacy information includes any opinions, concerns, preferences, activities, intentions, or any other information that could be considered private or sensitive based on the context of the query.
The phrase should be relevant to the user or the user’s related people or things based on the context of the query.
The phrase should not be a general term.
You should think step-by-step.
Finally, output your answer in JSON. If there is no phrase extracted, the output JSON list should be empty.
Here are some examples:
<|CASE|>End of examples.
The user’s current Query is
"<|QUERY|>"
Prompts for in-context learning (Chinese).
[…]
{CJK}UTF8gbsn 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
{CJK}UTF8gbsn 隐私短语应当有具体的指代,而不是一个泛指的词汇。
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
{CJK}UTF8gbsn 以下是一些例子:
{CJK}UTF8gbsn <|CASE|>例子结束。
{CJK}UTF8gbsn 用户当前的请求是:
"<|QUERY|>"
Table 24: Prompt template for in-context learning of privacy information summarization task in English and Chinese dataset, where <|INPUT|> is the input of user query and <|CASE|>is structured examples from training set.
Prompts for in-context learning (English).
[…]
Extract informative phrases from the query that reveal the privacy information of the user, or user related people or things.
The privacy information includes any opinions, concerns, preferences, activities, intentions, or any other information that could
be considered private or sensitive based on the context of the query.
The phrase should be relevant to the user or the user’s related people or things based on the context of the query.
Then for each extracted phrase, summarize the specific privacy information the phrase has leaked.
You should think step-by-step.
Finally, output your answer in JSON. If there is no phrase extracted, the output JSON list should be empty.
Here are some examples:
<|CASE|>End of examples.
The user’s current Query is
"<|QUERY|>"
Prompts for in-context learning (Chinese).
[…]
{CJK}UTF8gbsn 提取用户请求中,揭示了用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 隐私敏感信息包括用户的观点、偏好、活动、意图或根据请求上下文可能被认为是私人或敏感的信息。
{CJK}UTF8gbsn 提取的短语应当在请求的上下文中体现出与用户或用户相关的人员或事物的相关性。
{CJK}UTF8gbsn 然后,对于每个提取的短语,总结该短语泄露的具体隐私信息
{CJK}UTF8gbsn 请一步一步思考(Think step-by-step)。
{CJK}UTF8gbsn 给出理由并在最后,以JSON格式输出答案。如果没有提取到任何短语,则输出的JSON列表应为空。
{CJK}UTF8gbsn 以下是一些例子:
{CJK}UTF8gbsn <|CASE|>例子结束。
{CJK}UTF8gbsn 用户当前的请求是:
"<|QUERY|>"
Table 25: Instructions for training models for privacy leakage classification task, where <|INPUT|> is the input of user query.
Instruction for English datasets.
[…]
System:
Judge whether the query that reveal any privacy-sensitive information of the user, or user related people or things.
Message:
Query: "<|INPUT|>"
Instruction for Chinese datasets.
[…]
System:
{CJK}UTF8gbsn 判断Query是否泄漏了用户、用户相关人员或事物的隐私敏感信息
Message:
{CJK}UTF8gbsn Query: "<|INPUT|>"
Table 26: Instructions for training models for privacy phrase extraction task, where <|INPUT|> is the input of user query.
Instruction for English datasets.
[…]
System:
Extract informative phrases from the query that reveal the privacy-sensitive information of the user, or user related people or things.
Message:
Query: "<|INPUT|>"
Instruction for Chinese datasets.
[…]
System:
{CJK}UTF8gbsn 从Query中提取出能够揭示用户、用户相关人员或事物的隐私敏感信息的短语
Message:
{CJK}UTF8gbsn Query: "<|INPUT|>"
Table 27: Instructions for training models for privacy information summarization task, where <|INPUT|> is the input of user query.
Instruction for English datasets.
[…]
System:
Extract informative phrases from the query that reveal the privacy-sensitive information of the user, or user related people or things.
Then for each extracted phrase, summarize the specific privacy information the phrase has leaked.
Message:
Query: "<|INPUT|>"
Instruction for Chinese datasets.
[…]
System:
{CJK}UTF8gbsn 从Query中提取出能够揭示用户、用户相关人员或事物的隐私敏感信息的短语。
{CJK}UTF8gbsn 然后,针对每个提取的短语,总结该短语泄露的具体隐私信息
Message:
Query: "<|INPUT|>"